The OMX has been serviced , the OTFs have been updated and saved as default for SIM reconstructions. A 10x objective is available now to perform tile scans for sample overviews. The 4th camera is non functional due to a power failure, repair parts are on order. So for now, the far red detection is not possible.
A newly developed joystick is installed on the OMX-V4, it allows navigating through a sample, mark and start acquisition/mosaic. It is a bit fast but with practice will help you find the best cells more efficiently.
Few new features are implemented during the last service visit (for example: laser power tuning), ask CCI staff to inform you before working on the machine. Again, be sure to keep the OTFs of April 2018 for the reconstructions.
The OMX system has been serviced and new OTFs have been generated. The current OTF set that is selected in SoftWoRx should be used for reconstructions, do not change it!
The Bio-Imaging Utrecht community organizes an ImageJ/FIJI workshop for experienced users on Tuesday the 4th of April. Previous sessions were very well appreciated and they allow you to interact with experts in IJ coding. You can learn image analyses automation and the experts can help you find, tweak and apply the proper plugin(s) for you specific question. It is highly advised for microscopists who use FIJI but have problems overseeing the large analyses possibilities offered by the platform. Register using the form attached here (deadline 17th March)
On Tuesday 28th of February to Wednesday the 1st of March the OMX will be unavailable. A sCMOS PCO camera will be replaced and the system will be fully re-aligned.
The OMX has been serviced and new OTFs have been made. We recommend using these. A strange bug in the acquisition software causes the 405 nm laser line (1st channel) to present wrong values, i.e. intensity % of the laser in the drop-down list do not represent the actual values. The desktop background picture of the PC instructs you restore the right interface (in 3 clicks).
This week the filter drawer changer that was locked has been repaired. The system was serviced and softWoRx packages on the Windows master acquisition station and the workstations have been updated.
On the OMX imaging system, channel alignment is repeatedly disturbed by incorrect placement of the automated filter drawer (which sets in every time at startup) . A GE scientist is coming to solve this issue (in the week of February the 29th). Until then the filter drawer will be fixed to the BGR position to maintain correct aligning settings. We can work with it to image many but not all of your favourite dyes. Please take this into consideration.
The broken master PC on the OMX V4 system has been replaced and the OMX is fully functional again.
One of the PCs of the OMX-V4 system has a corrupted hard drive and does not boot. The system is not usable and will be repaired as soon as possible. The PC does not seem to be in stock in Europe, so it has been ordered in the US. As a result the system will not be operative for at least a week and likely for the rest if this month. We will do our best to speed up the process and will keep you posted in this blog.